OpenStudy (anonymous):

Help anyone, please? Rehan has been awarded some money in a settlement. He has the option to take a lump sum payment of $170,000 or get paid an annuity of $1,000 per month for the next 20 years. Which is the better deal for Rehan, and by how much, assuming the growth rate of the economy is 3.05% per year? Lump Sum: by $63,712.66 Lump Sum: by $7,707.19 Annuity: by $63,712.66 Annuity: by $7,707.19

OpenStudy (anonymous):

i think it's the 1st one?

OpenStudy (anonymous):

could you show me step by step?

OpenStudy (anonymous):

@Kainui help please?

OpenStudy (mathstudent55):

You need the formula for the present value of an annuity.

OpenStudy (mathstudent55):

\( P = A \times \dfrac{1 - (1 + i)^{-n}}{i} \) where P = present value A = annualized payment (per period) i = interest rate per period expressed as a decimal n = number of periods \( P = $1,000 \times \dfrac{1 - (1 + \frac{0.0305}{12})^{-{240}}}{\frac{0.0305}{12}} \) \(P = $179,499.73\) The present value of the annuity is higher than $170,000, but not by the amount in your choices. Is all the given info correct in your problem?

OpenStudy (anonymous):

Yes its all correct :/

OpenStudy (mathstudent55):

The difference is $9,499.73, not $7,707.19. I don't understand how your choices have that amount.

OpenStudy (anonymous):

@mathstudent55 :(